# Improving Quantum Approximate Optimization by Noise-Directed Adaptive Remapping

### v1

@@ Abstract
We present Noise-Directed Adaptive Remapping (NDAR), a heuristic meta-algorithm for approximately solving binary optimization problems by leveraging certain types of noise. We consider access to a noisy quantum processor with dynamics that features a global attractor state. In a standard setting, such noise can be detrimental to the quantum optimization performance. In NDAR, the algorithm bootstraps the attractor by iteratively gauge-transforming the cost-function Hamiltonian. In each iteration step, the gauge transformation effectively changes the attractor state into a higher-quality solution of the cost Hamiltonian based on the results of variational optimization in the previous step. The end result is that noise aids variational optimization, as opposed to hindering it. We demonstrate the effectiveness of our protocol in Quantum Approximate Optimization Algorithm (QAOA) experiments with subsystems of the newest generation of Rigetti Computing’s superconducting device Ankaa-2. We obtain approximation ratios (of best-found solutions) $0.9$-$0.96$ (spread across instances) for multiple instances of random, fully connected graphs (Sherrington-Kirkpatrick model) on $n=82$ qubits, using only depth $p=1$ (noisy) QAOA in conjunction with NDAR. This compares to $0.34$-$0.51$ for vanilla $p=1$ QAOA with the same number of function calls.
@@ I Introduction
Noise in state-of-the-art quantum processors limits quantum circuit depth. Significant effort has been spent to improve outer-loop optimization to variationally tune noisy quantum circuits [1], and to develop various error mitigation [2] and error correction [3, 4] techniques. In this work, we propose an algorithmic framework that does not mitigate the noise directly, but instead attempts to align the quantum optimization with the dissipative part of the evolution by using the knowledge about the noise.
@@
We refer to the new algorithm as Noise-Directed Adaptive Remapping (NDAR). The actual variational optimization method is a subroutine of NDAR that can be chosen in a way best suited for a given experimental setup (in fact, it does not necessarily have to be quantum). In this work, we focus mainly on the canonical example of Quantum Approximate Optimization Algorithm (QAOA) [5, 6] though the adaptive remapping approach we introduce applies to other approaches. We implement the NDAR adaptive loop for $p=1$ QAOA targeting the minimization of fully-connected, randomly-weighted Sherrington-Kirkpatrick Hamiltonians on $n=82$-qubit subsystem of Rigetti Computing’s quantum processing unit (QPU) based on superconducting transmons. We observe solutions with approximation ratios above $\approx 0.9$, compared to $\approx 0.5$ of vanilla QAOA, when both methods are given the same number of circuit runs.
@@ IV Experimental demonstration
We implement $p=1$ QAOA optimization for $n=82$ qubits with Tree of Parzen Estimators (TPE) [41, 42] parameter setting strategy using a budget of $t=100$ cost function evaluations (trials). Each trial corresponds to estimating the cost function value from $s=1000$ samples at fixed parameters. The QAOA is implemented for a set of 10 distinct Hamiltonian cost functions. For each Hamiltonian instance, the optimization is run for $20$ random bitflip gauges $P_{\mathbf{y}}$.
@@
The results are presented on the right side of Fig. 2. The data points correspond to the best-found solution, meaning the best bitstring out of $150*1000*(i+1)$ samples ($i$ being iteration step counting from 00). The NDAR QAOA results from QPU are colored orange. We also implement the vanilla QAOA (light blue) with the same number of function calls as NDAR. For reference, we provide the best approximation ratios obtained from the same number of shots using uniformly random sampling (grey).
@@ V Conclusions and future work
We introduced Noise-Directed Adaptive Remapping (NDAR), a new algorithmic framework for quantum approximate optimization in the presence of certain types of noise. NDAR iteratively changes the encoding of the target Hamiltonian in a way that adaptively attempts to align the quantum evolution with the noise dynamics. We believe that the Noise-Directed Adaptive Remapping algorithm has the potential to become a standard method in the general toolbox of quantum optimization in the non-fault-tolerant regime, and, more generally, a technique to exploit the interplay of symmetries and open-system dynamics to improve algorithm performance. In particular, as many superconducting, photonic, and electronic quantum information processing platforms are affected by amplitude-damping-like noise, we expect NDAR to be particularly useful for optimization implemented in those devices (including quantum annealers and analog simulators). In the previous sections, we empirically demonstrated the effectiveness of NDAR in a proof-of-concept $p=1$ vanilla QAOA optimization on Rigetti’s $82$-qubit QPU. However, more analytical and numerical work is required to better understand the general applicability of our method, as well as its dependence on the noise type and strength. Moreover, our study unveils numerous avenues for further research, a few of which we now outline.
First, we note that the proposed vanilla formulation of NDAR is intentionally very basic, and many heuristics can be applied to extend it. For example, the outer NDAR loop is currently greedy, meaning we always choose the new bitflip gauge based on the best-found solution. This has clear downsides, as follows from studies of classical greedy algorithms. It is possible that more flexible strategies would be beneficial – such as allowing for optimization over multiple low-energy gauges, performing the choice of the bitflip gauge based on its Hamming weight or typicality, or performing an additional local (classical) search in the neighborhood of the best-found solutions.
@@
Third, we believe it would be interesting to implement NDAR with different algorithms than QAOA. For example, NDAR might be conceptually related to recently studied dissipative quantum algorithms that aim to mimic the cooling processes in Nature to find low-energy states, see Refs. [47, 48, 49, 50, 51]. Bath-engineering techniques, which have shown promise to avoid the barren plateau problem, might also be coupled with our remappings [52]. Moreover, experimental tests with different types of unitary ansätze (including those that are not necessarily bitflip-gauge invariant, such as Quantum Mixer-Phaser Ansatz [53]) and optimization algorithms are warranted [6, 28, 54, 55, 56, 57, 1, 58, 59]. In particular, it would be intriguing to combine NDAR with methods such as ADAPT-QAOA [28]; with iterative/recursive versions of QAOA [60, 61, 62]; with warm-start techniques [29, 30, 31, 32] or with recently-proposed quantum optimization schemes that generate candidate solutions via expectation-values statistics [63, 64].
@@ Appendix B Numerical studies
Due to the above observations, we decided to highly reduce the number of trials and samples for the NDAR implementation – otherwise the NDAR would often converge at the first iteration (as indicated by results in Fig 4a), rendering the results uninteresting. The plot on the right of Fig. 4 shows the performance overview of the Noise-Directed Adaptive Remapping algorithm applied to noisy simulated QAOA. The QAOA is guided by the TPE optimizer with $t=20$ cost function evaluations, each involving gathering $s=100$ samples. We observe that the NDAR applied to noisy QAOA (orange) finds the ground state in all tested cases after 3 iterations. Results of vanilla QAOA (light blue) show that in that case, the performance is spread out across Hamiltonians, indicating that for some instances, the strong amplitude damping noise prohibits finding the ground state. The reference random uniform baseline (grey), given the same small number of samples, achieves comparable performance. The above results indicate that NDAR can aid the noisy QAOA even when the computational resources (function calls and samples) are highly limited.
@@ C.1 QAOA circuit implementation
The Rigetti’s Ankaa-2 (Fig. 5) device’s native gate set consists of iSWAP, ±$\pm\frac{\pi}{2}$ RX rotations and arbitrary RZ rotations. The device’s connectivity (w.r.t. iSWAP gate) is that of a rectangular 7 by 12 grid (with some edges missing, not shown in Fig. 5). To implement QAOA circuits for fully connected Hamiltonians, we implement an efficient SWAP network from Ref. [65] on linear chains of qubits. Since the phase separator (the $\exp\left(-i\gamma H\right)$ operator) of the QAOA ansatz implemented with SWAP network has a structure of multiple RZZ rotations followed by SWAP gates, we compile the product of the two gates jointly (a similar trick was applied for efficient compilation in Ref. [1]). Each implementation of RZZ+SWAP requires applying 3 ISWAP gates and at most 4 RX(±$\left(\pm\frac{\pi}{2}\right)$ and 5 RZ rotations (the exact number depends on the qubit). Implementing a single layer of QAOA ansatz for $n$ qubits requires consecutive application of $\approx\frac{n}{2}$ linear chains. Thus, for the biggest experiments shown in the text with $n=82$, the $p=1$ QAOA ansatz consists of around $\approx 10,000$ ISWAP gates and $\approx 30,000$ single-qubit rotations.
@@ C.2 Parameter setting strategy details
To implement variational optimization (for Figs. 1, 2, and 3) we use the Tree of Parzen Estimators (TPE) implementation from the Optuna package [42]. We use default hyperparameters (but with turned-off pruning trials). Each optimization was run with a random seed, and each optimization started from the same initial points (all angles set to $0.1$).

### v2

@@ Abstract
We present Noise-Directed Adaptive Remapping (NDAR), a heuristic algorithm for approximately solving binary optimization problems by leveraging certain types of noise. We consider access to a noisy quantum processor with dynamics that features a global attractor state. In a standard setting, such noise can be detrimental to the quantum optimization performance. Our algorithm bootstraps the noise attractor state by iteratively gauge-transforming the cost-function Hamiltonian in a way that transforms the noise attractor into higher-quality solutions. The transformation effectively changes the attractor into a higher-quality solution of the Hamiltonian based on the results of the previous step. The end result is that noise aids variational optimization, as opposed to hindering it. We present an improved Quantum Approximate Optimization Algorithm (QAOA) runs in experiments on Rigetti’s quantum device. We report approximation ratios $0.9$-$0.96$ for random, fully connected graphs on $n=82$ qubits, using only depth $p=1$ QAOA with NDAR. This compares to $0.34$-$0.51$ for standard $p=1$ QAOA with the same number of function calls.
@@ I Introduction
Noise in state-of-the-art quantum processors limits quantum circuit depth. Significant effort has been spent on improving the implementation of quantum approximate optimization [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], and developing various error mitigation [15] and error correction [16, 17] techniques. In this work, we propose an algorithmic framework that does not mitigate the noise directly, but instead attempts to align the quantum optimization with the dissipative part of the evolution by using the knowledge about the noise.
@@
We refer to the new algorithm as Noise-Directed Adaptive Remapping (NDAR). The actual variational optimization method is a subroutine of NDAR that can be chosen in a way best suited for a given experimental setup (in fact, it does not necessarily have to be quantum). In this work, we focus mainly on the canonical example of Quantum Approximate Optimization Algorithm (QAOA) [1, 2] though the adaptive remapping approach we introduce applies to other approaches. We implement the NDAR adaptive loop for $p=1$ QAOA targeting the minimization of fully-connected, randomly-weighted Sherrington-Kirkpatrick Hamiltonians on $n=82$-qubit subsystem of Rigetti Computing’s quantum processing unit (QPU) based on superconducting transmons. We observe solutions with approximation ratios above $\approx 0.9$, compared to $\approx 0.5$ of standard QAOA, when both methods are given the same number of circuit runs.
@@ IV Experimental demonstration
We implement $p=1$ QAOA optimization for $n=82$ qubits with Tree-Structured Parzen Estimators (TPE) [51, 52] parameter setting strategy using a budget of $t=100$ cost function evaluations (trials). Each trial corresponds to estimating the cost function value from $s=1000$ samples at fixed parameters. The QAOA is implemented for a set of 10 distinct Hamiltonian cost functions. For each Hamiltonian instance, the optimization is run for $20$ random bitflip gauges $P_{\mathbf{y}}$.
@@
The results are presented on the right side of Fig. 2. The data points correspond to the best-found solution, meaning the best bitstring out of $150*1000*(i+1)$ samples ($i$ being iteration step counting from 00). The NDAR QAOA results from QPU are colored orange. We also implement the standard QAOA (light blue) with the same number of function calls as NDAR. For reference, we provide the best approximation ratios obtained from the same number of shots using uniformly random sampling (grey).
@@ V Conclusions and future work
We introduced Noise-Directed Adaptive Remapping (NDAR), a new algorithmic framework for quantum approximate optimization in the presence of certain types of noise. NDAR iteratively changes the encoding of the target Hamiltonian in a way that adaptively attempts to align the quantum evolution with the noise dynamics. We believe that the Noise-Directed Adaptive Remapping algorithm has the potential to become a standard method in the general toolbox of quantum optimization in the non-fault-tolerant regime, and, more generally, a technique to exploit the interplay of symmetries and open-system dynamics to improve algorithm performance. In particular, as many superconducting, photonic, and electronic quantum information processing platforms are affected by amplitude-damping-like noise, we expect NDAR to be particularly useful for optimization implemented in those devices (including quantum annealers and analog simulators). In the previous sections, we empirically demonstrated the effectiveness of NDAR in a proof-of-concept $p=1$ standard QAOA optimization on Rigetti’s $82$-qubit QPU. However, more analytical and numerical work is required to better understand the general applicability of our method, as well as its dependence on the noise type and strength. Moreover, our study unveils numerous avenues for further research, a few of which we now outline.
First, we note that the proposed formulation of NDAR is intentionally very basic, and many heuristics can be applied to extend it. For example, the outer NDAR loop is currently greedy, meaning we always choose the new bitflip gauge based on the best-found solution. This has clear downsides, as follows from studies of classical greedy algorithms. It is possible that more flexible strategies would be beneficial – such as allowing for optimization over multiple low-energy gauges, performing the choice of the bitflip gauge based on its Hamming weight or typicality, or performing an additional local (classical) search in the neighborhood of the best-found solutions.
@@
Third, we believe it would be interesting to implement NDAR with different algorithms than QAOA. For example, NDAR might be conceptually related to recently studied dissipative quantum algorithms that aim to mimic the cooling processes in Nature to find low-energy states, see Refs. [57, 58, 59, 60, 61]. Bath-engineering techniques, which have shown promise to avoid the barren plateau problem, might also be coupled with our remappings [62]. Moreover, experimental tests with different types of unitary ansätze (including those that are not necessarily bitflip-gauge invariant, such as Quantum Mixer-Phaser Ansatz [63]) and optimization algorithms are warranted [2, 6, 64, 65, 66, 67, 8, 68, 69, 70]. In particular, it would be intriguing to combine NDAR with methods such as ADAPT-QAOA [6]; with iterative/recursive versions of QAOA [71, 72, 73]; with warm-start techniques [39, 40, 41, 42] or with recently-proposed quantum optimization schemes that generate candidate solutions via expectation-values statistics [74, 75].
@@ Appendix B Numerical studies
Due to the above observations, we decided to highly reduce the number of trials and samples for the NDAR implementation – otherwise the NDAR would often converge at the first iteration (as indicated by results in Fig 4a), rendering the results uninteresting. The plot on the right of Fig. 4 shows the performance overview of the Noise-Directed Adaptive Remapping algorithm applied to noisy simulated QAOA. The QAOA is guided by the TPE optimizer with $t=20$ cost function evaluations, each involving gathering $s=100$ samples. We observe that the NDAR applied to noisy QAOA (orange) finds the ground state in all tested cases after 3 iterations. Results of standard QAOA (light blue) show that in that case, the performance is spread out across Hamiltonians, indicating that for some instances, the strong amplitude damping noise prohibits finding the ground state. The reference random uniform baseline (grey), given the same small number of samples, achieves comparable performance. The above results indicate that NDAR can aid the noisy QAOA even when the computational resources (function calls and samples) are highly limited.
@@ C.1 QAOA circuit implementation
The Rigetti’s Ankaa-2 (Fig. 5) device’s native gate set consists of iSWAP, ±$\pm\frac{\pi}{2}$ RX rotations and arbitrary RZ rotations. The device’s connectivity (w.r.t. iSWAP gate) is that of a rectangular 7 by 12 grid (with some edges missing, not shown in Fig. 5). To implement QAOA circuits for fully connected Hamiltonians, we implement an efficient SWAP network from Ref. [76] on linear chains of qubits. Since the phase separator (the $\exp\left(-i\gamma H\right)$ operator) of the QAOA ansatz implemented with SWAP network has a structure of multiple RZZ rotations followed by SWAP gates, we compile the product of the two gates jointly (a similar trick was applied for efficient compilation in Ref. [8]). Each implementation of RZZ+SWAP requires applying 3 ISWAP gates and at most 4 RX(±$\left(\pm\frac{\pi}{2}\right)$ and 5 RZ rotations (the exact number depends on the qubit). Implementing a single layer of QAOA ansatz for $n$ qubits requires consecutive application of $n$ linear chains, each with $\approx\frac{n}{2}$ RZZ+SWAP gates. Thus, for the biggest experiments shown in the text with $n=82$, the $p=1$ QAOA ansatz consists of around $\approx 10,000$ ISWAP gates and $\approx 30,000$ single-qubit rotations.
@@ C.2 Parameter setting strategy details
To implement variational optimization (for Figs. 1, 2, and 3) we use the Tree-Structured Parzen Estimators (TPE) implementation from the Optuna package [52]. We use default hyperparameters (but with turned-off pruning trials). Each optimization was run with a random seed, and each optimization started from the same initial points (all angles set to $0.1$).